# Explain to Question not to Justify

### v1

@@ 4 Discussion
Start explaining models for high stakes decisions. The need for explainability is often questioned and criticized. One of the most well-known arguments against explaining complex models are presented in the paper ,,Stop explaining black box machine learning models for high stakes decisions and use interpretable models instead”Rudin (2019). Note, however, that the criticism presented there relates to the objectives indicated with BLUE XAI. Under this general title, the author criticizes the use of models that may not work, but their behavior is masked or justified by XAI-washing. The objectives in RED XAI are completely different and the RED XAI process itself can and should be used for any type of model, especially models used in high-stake decisions. While it is certainly true that providing explanations to the lay user can lead to an unsubstantiated trust in the system Weller (2019), we strongly believe that an increased understanding of the functioning of the model and a RED team mentality are necessary ingredients to build truly reliable AI systems in the future.
Why two cultures and not five or ten? Many of the review papers discussed in the second section identified a number of XAI stakeholders. While such fragmentation often allows us to capture all or almost all potential uses of XAI, it does not create a mental mindset that is easy to remember and then apply. Therefore, in this paper we have intentionally reduced the multitude of XAI applications by reducing them to two archetypal cultures. We are fully aware that there will be applications that do not fit into the two sketched poles. This, however, is fine. Untangling the research threads in XAI community into the two described cultures clarifies thinking about new research directions, and this clarification is what we aimed for.
Frontier models (aka foundation models), unlike classical predictive models, can be applied to a wide variety of tasks. Typically, new applications of such models were not considered in the process of training, and often they were not even known, and only the availability of the model made it possible to see new applications. In such situations, model exploration and debugging techniques are even more essential. The more potential applications, the greater the potential safety problem.

### v2

@@ 4 Discussion
In this section, we include topics pointed out during the review process. Readers are encouraged to expand or add more topics to the two XAI cultures.
Figure 4: Illustration of questions and needs raised by stakeholders interested in Blue XAI and Red XAI techniques. Questions are framed using the example of a hypothetical covid death risk assessment model used to prioritize access to vaccination. For each of the seven example areas, a sample question is presented that can be answered using XAI techniques.
@@
Start explaining models for high-stakes decisions. The need for explainability is often questioned and criticized. One of the most well-known arguments against explaining complex models are presented in the paper “Stop explaining black box machine learning models for high stakes decisions and use interpretable models instead”(Rudin, 2019). Note, however, that the criticism presented there relates to the objectives indicated with BLUE XAI. Under this general title, the author criticizes the use of models that may not work, but their behavior is masked or justified by XAI-washing. The objectives in RED XAI are completely different and the RED XAI process itself can and should be used for any type of model, especially models used in high-stakes decisions. While it is certainly true that providing explanations to the lay user can lead to an unsubstantiated trust in the system (Weller, 2019), we strongly believe that an increased understanding of the functioning of the model and a RED team mentality are necessary ingredients to build truly reliable AI systems in the future.
Frontier models (aka foundation models), unlike classical predictive models, can be applied to a wide variety of tasks. Typically, new applications of such models were not considered in the process of training, and often they were not even known, and only the availability of the model made it possible to see new applications. In such situations, model exploration and debugging techniques are even more essential. The more potential applications, the greater the potential safety problem. The BLUE-XAI community may approach the topic of LLM/foundation models from a “human alignment” perspective. It expects models to respect human values and not discriminate, cheat, use toxic speech, or hate speech. Benchmarks are being created in these directions, verifying how truthful, ethical, and “aligned” the LLMs/foundation models are. The RED-XAI community may approach this problem from a different angle. Understanding the model is needed to control the model. One example of such an application is an experiment with the editing property of latent representation (Meng et al., 2022), to better understand how LLMs store facts. Such knowledge opens new opportunities, e.g. to force the model to forget the facts in question. Another approach is the study of grooking phenomenon (Nanda et al., 2023), whose goal is to understand the more complex models through studying simpler/smaller models.
Covid-19 use-case shown in Figure 4 illustrates the differences and similarities between the Blue and Red XAI perspectives. The model for assessing the risk of death could be used in prioritizing access to certain medical interventions, such as access to vaccines. The Blue XAI part is focused on the perspective of the user, in the case of this model it could be the perspective of a patient or a doctor. The transparency of such a model is important to achieve a high level of trust from users. The Red XAI part is focused on what information can be extracted from the model (which is important in the case of a new disease about which little was known) and how the model can be more accurately tested and fixed.
Purple XAI The terminology of the Red Team and Blue Team is rooted in military and security. Over time, additional ‘colors’ have been introduced into the team catalog, and this may also be the case with XAI communities. For example, Purple XAI can denote methods that can be used by both the model developer and the end users (certainly the most popular SHAP methods fall into this group). Similarly, using the color scheme used in security, one might discuss Yellow XAI which is more focused on the creation of processes, frameworks, and benchmarks that evaluate XAI methods, without explicitly defining an audience. However, no matter how many different subgroups of the XAI community can be distinguished, it is worthwhile for each of these communities to have clear objectives.
Secret sauce The methods developed by Blue XAI are targeted at the end user and will therefore be publicly visible quickly. The nature of the methods used by Red XAI may be different, as the tools that lead to better models may be regarded as a competitive advantage and protected know-how that model-creating organizations may not want to share (whether for fear of model security or competition). It is therefore possible that some of the new Red XAI methods will be developed outside the public academic community, and this may also explain the lower visibility of such methods at present.
Acknowledgments
PB acknowledges financial support from the SONATA BIS grant 2019/34/E/ST6/00052 funded by Polish National Science Centre (NCN). WS acknowledges financial support by the German Research Foundation (DFG) - Research Unit KI-FOR 5363 (project ID: 459422098).
Impact Statement
This paper aims at advancing the field of Trustworthy Machine Learning. The work has no immediate social implications, although in the long run we believe that sorting out the threads related to explainability will lead to the development of better models.